Can I please get some help with this problem.
Solve the following system of two linear equations.
5x-6y= 40
3x+2y=8
Thank you

Answer by mananth(16946)

5x-6y= 40................1
3x+2y=8...................2
...
multiply equation (2) by 3
9x+6y=24
...
add it to (1)
9x+5x+6y-6y=24+40
14x=64
/14
x=64/14
x=32/7
...
plug value of x in (1)
5*(32/7)-6y=40
6y=5(32/7)-40
y=-20/7
..
m.ananth@hotmail.ca